The lengths of a professor's classes has a continuous uniform distribution between 50.0 min and 52.0 min. If one such class is randomly selected, find the probability that the class length is more than 51.6 min. P(X > 51.6) =
Answer #1

Solution :

uniform distribution properties and rules

Given that,

a = 50

b = 52

P(x > c) = (b - c) / (b - a)

P(x > 51.6) = (52 - 51.6) / (52 - 50) = 0.2

Probability = 0.2